2. Find all integers x #= 3 such that x-3Ix3-3.


1
Expert's answer
2022-06-27T09:34:36-0400

"\dfrac {x^3-3}{x-3}=~ \dfrac {x^3-27+24}{x-3}= ~\dfrac {(x-3)(x²+3x+9)+24}{x-3}=~\underbrace{x²+3x+9}_{\text{integer}}~+~\dfrac {24}{x-3}"


So, we are looking for values of x, such that:


"x-3 ~|~24 \implies \\
\implies ~x-~3\in~\begin{Bmatrix}
 ±1,±2,±3,±4,±6,±8,±12,±24
\end{Bmatrix} \implies \\
\implies ~x \in ~ \begin{Bmatrix}
−21,−9,−5,-3,−1,0,1,2,4,5,6,7,9,11,15,27
\end{Bmatrix}"



Corresponding x values are


"\u221221,\u22129,\u22125,\u22123,\u22121,0,1,2,4,5,6,7,9,11,15 ~and~ 27"